# Is Hyper-Parameter Optimization Different for Software Analytics?

### v3

@@ 1 Introduction
This is an important result since HPO is an important problem. Hyper-parameter optimizers explore the myriad of configuration choices which control either the learning or the data pre-processing. As shown by Table I and Table II, there are many such hyper-parameters. Much SE research has shown that HPO is very effective at improving the performance of models built from SE data [24, 2, 70, 72]; for example, see Figure 1. Nevertheless, while studying industrial text-mining tools, we have found many examples of poorly-chosen hyper-parameters (which meant our industrial partners were selling less-than-optimal products) [41].
@@
SE data can be different to other kinds of data; and those differences mean that we should use different kinds of algorithms for our data.
The rest of this paper is structured as follows. The next section discusses how SE data is different to AI data. This is followed by notes on related work, specifically, hyperparameter optimization and its role in software analytics. Section 5 discusses our experimental setup, and the results of those experiments are presented in Section 6. We discuss why SMOOTHIE works so well for SE data along with threats to validity in Section 7, before discussing limitations and concluding.
@@ 2.1 Expect Smoothness in SE Data
To say that another way, a learner that respects the smoothness of SE data does better than otherwise. Based on that observation, we theorized that:
@@ 3 Hyper-parameter Optimization (HPO)
TABLE I: A sample hyper-parameter choices for learners (note, this is not a complete set).
@@
TABLE II: A sample of hyper-parameter choices for data pre-processing (note, this is not a complete set).
@@ 3.1 About HPO
A learning system combines some data pre-processing and machine learning. Every such system has magic control parameters that adjust the data pre-processing and/or the machine learning (e.g. see Table I and Table II). Even something as simple as the $k-$nearest neighbors classifier needs to decide how many neighbors it should study in order to classify a new example (and, as seen in Table I, it must settle on several other configuration options).
@@ 3.2 HPO in software analytics
As shown in Figure 1, many of the algorithms used in defect prediction can benefit from hyper-parameter optimization. Recently, Yedida and Menzies [70] proposed “fuzzy sampling” (see Table II) and combined it with stacked autoencoders, SMOTE, hyper-parameter optimization (using DODGE [2]), and a weighted loss function to achieve state-of-the-art results on defect prediction using the PROMISE repository [10], which contains 20 static code attributes of the software such as the depth of inheritance tree [15], and mean McCabe cyclomatic complexity [51]. In this study, we use the same 6 datasets as that study for a fair comparison.
@@ 4 About SMOOTHIE
SMOOTHIE begins by sampling a few configurations of a learner, selected at random from, say, Table II. Then, we compute the smoothness for each configuration using the methods described in this section.
@@ 8       P[config] $\leftarrow$ Run(config);
  • In some cases smoothness can be computed just by looking at certain properties of the data (in Gaussian Naive Bayes, that is the covariance matrix of the data, see Theorem 2). In this first case, HPO is very fast–just peek at the data to rank and prune the operators of (say) Table II.
@@ 4.2 Smoothness for Feedforward Classifiers
This section shows the smoothness computation for different learners. We prove the result for feedforward classifiers here; for other derivations (Naive Bayes, logistic regression), please see the Supplementary Material.
@@
Note: It is typical to use a regularization term in addition to the standard loss function above. In the supplementary material, we show that the addition of the regularizer only adds a constant term to the smoothness, and as such, it can be disregarded.
@@ 4.4 Smoothness for Gaussian Naive Bayes
For the proof of the following, please see the supplementary material.
@@ 4.5 Smoothness for other learners
TABLE III: Dataset sizes for each of the tasks in this study. Numbers on brackets denote number of SE papers we could find that have explore these options (in a Google search of "hyper-parameter optimization" and "software engineering" 2014-2024).
@@ 5.1 Data
Table III describes the datasets used in this study. Clearly, our study encompasses a variety of data shapes: both small and large, and both wide and narrow. In particular, for static code warnings, the number of features varies by dataset because of the data collection process from FindBugs; see [38] for a detailed report of the data cleaning process.
@@ 5.3 Hyper-Parameter Options
Our hyper-parameter options come from Tables III, and are summarized in Figure 5. We use these options since, as detailed in Table II, all those options have found favor with prior researchers. In Figure 5, branches with a solid line represent an AND, while dashed lines represent an OR. The learner-independent space has 60 options, and the learners can be configured in 100 ways. This yields 60 * 100 = 6,$0$ options in total. For the experiments on the Bayesmark datasets, we use the hyper-parameter space used by Bayesmark, which for feedforward networks optimized by Adam, has a size of 134M options [^7] https://github.com/uber/bayesmark/blob/master/bayesmark/sklearn_funcs.py#L91-L101.
@@ 5.5 Evaluation Metrics
TABLE IV: RQ1 results: Static code warnings results. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). For Algorithm 1, we use $N_{1}=50,N_{2}=10$. FF = Feedforward network. NB = Naive Bayes. LR = Logistic Regression.
@@ 6 Results
TABLE V: RQ1 results: Issue lifetime prediction results. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). Our baseline is DeepTriage [49]. In the dataset column, the number of classes the data is binned into is in parentheses.
@@
TABLE VI: RQ1 results: Defect prediction results. For Algorithm 1, we use $N_{1},N_{2}$ =30, 5. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
TABLE VII: RQ1 results: Results on the NeurIPS 2020 Black-Box Optimization Challenge public datasets. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
@@ 6.1 RQ1: What are the benefits of smoothing?
Table IV shows our results on static code analysis false alarms. Recall that the task here was to recognize static code warnings that, in the historical log, were ignored by developers. In the table, gray denotes “best results in this column, for that dataset“. Note that SMOOTHIE (and more specifically, SMOOTHIE (FF) has much more gray than anything else. More specifically: all the non-SMOOTHIE methods have far more white cells (i.e. non best) than the SMOOTHIE methods. (Table VIII shows the exact win/tie counts).
Table V shows our results on issue lifetime prediction. Recall that the task here is to classify a GitHub issue according to how long it will take to resolve. As done in prior work [43], for 3-class classification, the number of days taken to fix bugs is binned into 0 to 1 days, 2 to 5 days, and 5 to 100 days. For 2-class, the number of days taken to fix bugs is binned into less than (or greater than) 4 days for Firefox, less than (or greater than) 5 days for Chromium, and less than (or greater than) 6 days for Eclipse. Note that:
  • For two class prediction (eclipse(2), firefox(2), chromium(2)) SMOOTHIE does not perform well (as shown in Table VIII, 1 tie and 2 losses). These datasets have the highest smoothness (i.e., sharpest minima) over all our SE datasets: the 2-class Chromium, Eclipse, and Firefox datasets had smoothness values of 9.4, 14.9, and 13.8 respectively. As discussed in RQ3 (Figure 7), we expect SMOOTHIE to perform poorly on datasets with such a high smoothness. This explains why SMOOTHIE did not work well for the 2-class binned version of the data.
@@
TABLE VIII: RQ1 results (summary): from 20 repeats over 24 data sets.
@@
Table VI shows defect prediction results. Recall that this task imports static code features and outputs a prediction of where bugs might be found. It turns out that defect prediction is our simplest task (evidence: all methods achieve the same best ranked results); defect prediction may not be a useful domain to compare optimizers.
Table VII shows our results for the NeurIPS 2020 Black- Box Optimization Challenge public dataset. We include these results here just to study how well SMOOTHIE performs on non-SE data.
Table VIII summarizes the RQ1 results from Table IV to Table VII. In that table, two methods tie if, they achieve the same statistical rank (using the methods of §5.6). Otherwise, a win is when one measure has a different rank to another and one measure has better median values (where “better” means “less“ for false alarms and “greater” for other metrics).
In summary, from Table VIII, we say:
@@
Figure 6: RQ2 Results: Runtime results on the chromium-2 class dataset.
@@ 6.3 RQ3: Should we use different kinds of algorithms for our SE data?
Apart from performance and runtime details, what does the above mean for the challenge problem sketched out in the introduction to this paper? To answer that question, we first recall Figure 3 where the mean smoothness for our SE and AI data has values 2 and 5, respectively.
Turning now to Figure 7, that figure reports the probability of SMOOTHIE outperforming BOHB, measured as the fraction of datasets for which SMOOTHIE outperforms BOHB (measured in terms of the F1 metric) [^8] More precisely, $\forall i>0,f(x_{i})=\frac{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\land(\text{SMOOTHIE}>\text{BOHB})\}}{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\}}$
@@ 7.1 Why Does SMOOTHIE Work?
As to why SMOOTHIE fails for bumpy spaces, we say its design trade-offs do not favor such a space, for the following reason. We conjecture that SMOOTHIE flattens the loss function by some ratio of, say, 50%. If this were true then when $\beta$-smoothness has a large value of, say, 5.6 (see Figure 9(d).d), SMOOTHIE can only flatten the landscape to .5*5.6=2.8 (see Figure 9(d).c, which is still quite steep). On the other hand, if we start near the mean smoothness of our SE data (see Figure 9(d).b) then a 50% flattening takes us into a quite flat region (see Figure 9(d).a) where models generalize better.
@@
Figure 9: An illustration of sharp minima forming with higher smoothness. Figure 9(d)d is the upper limit of the sharpness that SMOOTHIE can handle (see Figure 7).
All figures generated using Wolfram Alpha.
@@ 7.2 Threats to Validity
8 Limitations and Future Work
The major limitation of this work is the choice of measurement of smoothness, and the use of smoothness as a heuristic. For example, having derived the Hessian, the condition number is an alternate measure of the “nice nature” of the landscape [11]. Another clear limitation is that this approach is limited to learners for which a smoothness can be derived. Especially in the case where the Hessian may not exist (e.g., if the loss function is not twice-differentiable such as the mean absolute error loss), this approach fails. Nevertheless, we have shown successful application in a classical learner that is not gradient-based, Naive Bayes. In that case, we used the negative log-likelihood as a surrogate for the loss function. We postulate that this should be possible for other learners too, such as the dual formulation in SVMs or the negative log-likelihood in decision trees. This issue would a frutiful direction for future work. Finally, we reiterate that SMOOTHIE performs poorly on datasets with high smoothness, as evidenced by our results on the 2-class issue close time prediction results.

### v4

@@ 1 Introduction
This is an important result since HPO is an important problem. Hyper-parameter optimizers explore the myriad of configuration choices which control either the learning or the data pre-processing. As shown by Table II and Table III, there are many such hyper-parameters. Much SE research has shown that HPO is very effective at improving the performance of models built from SE data [24, 2, 74, 76]; for example, see Figure 1. Nevertheless, while studying industrial text-mining tools, we have found many examples of poorly-chosen hyper-parameters (which meant our industrial partners were selling less-than-optimal products) [43].
@@
SE data can be different from other kinds of data; and those differences mean that we should use different kinds of algorithms for our data.
To summarize, the core contributions of this paper are:
  • We extend empirical observations from prior work on static code warnings, which showed that SE loss landscapes can be “smoother” than non-SE landscapes, to show that this is a domain-specific idiosyncrasy of SE data. We connect this observation with intuitive explanations for why SE data exhibits this property, while standard AI data cannot. The crux of this paper is that this property is specific to SE data, and that SE research should devise algorithms that leverage such unique properties.
  • We exploit this property to devise a hyper-parameter optimization method tailored to SE data. Our method uses a very cheap probing computation to test the smoothness of loss landscapes, discarding configurations that are unpromising. We motivate this method with recent work from the AI literature showing that such “smooth” landscapes are desirable for generalization to unseen data.
  • We derive semi-empirical equations to compute the smoothness in a computationally cheap manner for different learning algorithms. We then experiment on three SE tasks (for a total of 19 datasets), showing that our method outperforms standard AI methods, while also being faster. Further, we show that the assumption underpinning this method (that of “smooth” loss landscapes) does not hold in general for AI datasets–which means standard AI methods may be preferable in such scenarios.
However, it is important that the reader be aware of the limitations of our method (these are expanded upon in the Limitations section at the end of this paper):
  • As presented here, our method relies on deriving learner-specific equations for smoothness. In future work, we will extend this to more learning algorithms and also devise a learner-agnostic method; but that is out of scope of this paper.
  • The above limitation leads to some complications in implementations. We provide a reference implementation for our method, which uses function overloading to use the type of the classifier passed to select the corresponding equation.
The rest of this paper is structured as follows. The next section discusses how SE data is different to AI data. This is followed by notes on related work, specifically, hyperparameter optimization and its role in software analytics. Section 5 discusses our experimental setup, and the results of those experiments are presented in Section 6. We discuss why SMOOTHIE works so well for SE data along with threats to validity in Section 7, before discussing limitations in Section 8 and future work in Section 9.
@@ 2.1 Expect Smoothness in SE Data
To say that another way, a learner that respects the smoothness landscape of SE data does better than otherwise.
TABLE I: Different kinds of landscapes
Loss landscape: Loss landscapes are the concept this paper predominantly talks about, and refers to the topography of the loss function of a deep learner. This is a function of the model’s parameters, and therefore exists in a high-dimensional space.
Hyper-parameter landscape: The hyper-parameter landscape [58, 35] is the topography of a loss function whose inputs are the hyper-parameters chosen. This function typically involves the model’s loss function, and in an empirical risk minimization formulation can be written as a bilevel optimization problem
$h^{*}&=\operatorname*{arg\,max}\limits_{h\sim\mathcal{H}}\mathbb{E}_{\mathcal{X\sim D}}\left[f(\theta^{*}(h))\right]\\ \theta^{*}(h)&=\operatorname*{arg\,min}\limits_{\theta\in\mathbb{R}^{d}}\mathcal{L}(X;\theta,h)\end{aligned}$ where $f$ is a performance measure (such as accuracy), and for brevity, we have omitted that the model is typically evaluated on a different dataset held out from the training set. While the hyper-parameter optimization problem itself is over this landscape, we use a random search to explore it in this paper, and for the chosen hyper-parameters at each step, we use the smoothness to decide if it is worth exploring further.
(Aside: it should be noted that the machine learning literature distinguishes “loss landscapes” from the “hyperparameter landscape”, which is the loss value distribution over the space of hyper-parameters). So, to be clear, when we say “landscape” we are referring to the loss landscape (for more on this, please see Table I).
Based on that observation, we theorized that:
@@ 3 Hyper-parameter Optimization (HPO)
TABLE II: A sample hyper-parameter choices for learners (note, this is not a complete set).
@@
TABLE III: A sample of hyper-parameter choices for data pre-processing (note, this is not a complete set).
@@ 3.1 About HPO
A learning system combines some data pre-processing and machine learning. Every such system has magic control parameters that adjust the data pre-processing and/or the machine learning (e.g. see Table II and Table III). Even something as simple as the $k-$nearest neighbors classifier needs to decide how many neighbors it should study in order to classify a new example (and, as seen in Table II, it must settle on several other configuration options).
@@ 3.2 HPO in software analytics
As shown in Figure 1, many of the algorithms used in defect prediction can benefit from hyper-parameter optimization. Recently, Yedida and Menzies [74] proposed “fuzzy sampling” (see Table III) and combined it with stacked autoencoders, SMOTE, hyper-parameter optimization (using DODGE [2]), and a weighted loss function to achieve state-of-the-art results on defect prediction using the PROMISE repository [10], which contains 20 static code attributes of the software such as the depth of inheritance tree [15], and mean McCabe cyclomatic complexity [53]. In this study, we use the same 6 datasets as that study for a fair comparison.
@@ 4 About SMOOTHIE
SMOOTHIE begins by sampling a few configurations of a learner, selected at random from, say, Table III. Then, we compute the smoothness for each configuration using the methods described in this section.
@@ 8       P[config] $\leftarrow$ Run(config);
  • In some cases smoothness can be computed just by looking at certain properties of the data (in Gaussian Naive Bayes, that is the covariance matrix of the data, see Theorem 2). In this first case, HPO is very fast–just peek at the data to rank and prune the operators of (say) Table III.
@@ 4.1 Preliminaries
Motivated by the unique properties of landscapes generated by SE data, SMOOTHIE solves the following problem:
$\operatorname*{arg\,max}\limits_{h\in\mathcal{H}}\max\limits_{x\sim\mathcal{X}}\left\lVert\nabla^{2}_{W^{[L]}}\mathcal{L}(f;x,h)\right\rVert$
Notice that choosing a specific hyper-parameter configuration and dataset materializes a specific loss function, whose smoothness we can compute. We are instead interested in a bound on the smoothness of all possible landscapes given a hyper-parameter configuration, which the inner $\max$ computes. The outer $\max$ searches across the hyper-parameter space itself in a randomized fashion.
@@ 4.2 Smoothness for Feedforward Classifiers
This section shows the smoothness computation for different learners. We prove the result for feedforward classifiers here; for other derivations (Naive Bayes, logistic regression), please see the proofs at the end of this paper.
@@
Note: It is typical to use a regularization term in addition to the standard loss function above. In the proofs section (at the end of this paper), we show that the addition of the regularizer only adds a constant term to the smoothness, and as such, it can be disregarded.
@@ 4.4 Smoothness for Gaussian Naive Bayes
For the proof of the following, please see the proofs section at the end of this paper.
@@ 4.5 Smoothness for other learners
TABLE IV: Dataset sizes for each of the tasks in this study. Numbers on brackets denote number of SE papers we could find that have explore these options (in a Google search of "hyper-parameter optimization" and "software engineering" 2014-2024).
@@ 5.1 Data
Table IV describes the datasets used in this study. Clearly, our study encompasses a variety of data shapes: both small and large, and both wide and narrow. In particular, for static code warnings, the number of features varies by dataset because of the data collection process from FindBugs; see [40] for a detailed report of the data cleaning process.
@@ 5.3 Hyper-Parameter Options
Our hyper-parameter options come from Tables IIIII, and are summarized in Figure 5. We use these options since, as detailed in Table III, all those options have found favor with prior researchers. In Figure 5, branches with a solid line represent an AND, while dashed lines represent an OR. The learner-independent space has 60 options, and the learners can be configured in 100 ways. This yields 60 * 100 = 6,$0$ options in total. For the experiments on the Bayesmark datasets, we use the hyper-parameter space used by Bayesmark, which for feedforward networks optimized by Adam, has a size of 134M options [^7] https://github.com/uber/bayesmark/blob/master/bayesmark/sklearn_funcs.py#L91-L101.
@@ 5.5 Evaluation Metrics
TABLE V: RQ1 results: Static code warnings results. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). For Algorithm 1, we use $N_{1}=50,N_{2}=10$. FF = Feedforward network. NB = Naive Bayes. LR = Logistic Regression.
@@ 6 Results
TABLE VI: RQ1 results: Issue lifetime prediction results. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). Our baseline is DeepTriage [51]. In the dataset column, the number of classes the data is binned into is in parentheses.
@@
TABLE VII: RQ1 results: Defect prediction results. For Algorithm 1, we use $N_{1},N_{2}$ =30, 5. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
TABLE VIII: RQ1 results: Results on the NeurIPS 2020 Black-Box Optimization Challenge public datasets. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
@@ 6.1 RQ1: What are the benefits of smoothing?
Table V shows our results on static code analysis false alarms. Recall that the task here was to recognize static code warnings that, in the historical log, were ignored by developers. In the table, gray denotes “best results in this column, for that dataset“. Note that SMOOTHIE (and more specifically, SMOOTHIE (FF) has much more gray than anything else. More specifically: all the non-SMOOTHIE methods have far more white cells (i.e. non best) than the SMOOTHIE methods. (Table IX shows the exact win/tie counts).
Table VI shows our results on issue lifetime prediction. Recall that the task here is to classify a GitHub issue according to how long it will take to resolve. As done in prior work [45], for 3-class classification, the number of days taken to fix bugs is binned into 0 to 1 days, 2 to 5 days, and 5 to 100 days. For 2-class, the number of days taken to fix bugs is binned into less than (or greater than) 4 days for Firefox, less than (or greater than) 5 days for Chromium, and less than (or greater than) 6 days for Eclipse. Note that:
  • For two class prediction (eclipse(2), firefox(2), chromium(2)) SMOOTHIE does not perform well (as shown in Table IX, 1 tie and 2 losses). These datasets have the highest smoothness (i.e., sharpest minima) over all our SE datasets: the 2-class Chromium, Eclipse, and Firefox datasets had smoothness values of 9.4, 14.9, and 13.8 respectively. As discussed in RQ3 (Figure 7), we expect SMOOTHIE to perform poorly on datasets with such a high smoothness. This explains why SMOOTHIE did not work well for the 2-class binned version of the data.
@@
TABLE IX: RQ1 results (summary): from 20 repeats over 24 data sets.
@@
Table VII shows defect prediction results. Recall that this task imports static code features and outputs a prediction of where bugs might be found. It turns out that defect prediction is our simplest task (evidence: all methods achieve the same best ranked results); defect prediction may not be a useful domain to compare optimizers.
Table VIII shows our results for the NeurIPS 2020 Black- Box Optimization Challenge public dataset. We include these results here just to study how well SMOOTHIE performs on non-SE data.
Table IX summarizes the RQ1 results from Table V to Table VIII. In that table, two methods tie if, they achieve the same statistical rank (using the methods of §5.6). Otherwise, a win is when one measure has a different rank to another and one measure has better median values (where “better” means “less“ for false alarms and “greater” for other metrics).
In summary, from Table IX, we say:
@@
Figure 6: RQ2 Results: Runtime on chromium-2 dataset.
@@ 6.3 RQ3: Should we use different kinds of algorithms for our SE data?
Turning now to the problem sketched in our introduction, recall Figure 3 where the mean smoothness of SE, AI data was 2,5, respectively. Figure 7, shows the reports the probability of SMOOTHIE outperforming BOHB’s F1 score, measured as the fraction of datasets where SMOOTHIE outperforms BOHB [^8] More precisely, $\forall i>0,f(x_{i})=\frac{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\land(\text{SMOOTHIE}>\text{BOHB})\}}{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\}}$
@@ 7.1 Why Does SMOOTHIE Work?
As to why SMOOTHIE fails for bumpy spaces, we say its design trade-offs do not favor such a space, for the following reason. We conjecture that SMOOTHIE flattens the loss function by some ratio of, say, 50%. If this were true then when $\beta$-smoothness has a large value of, say, 5.6 (see Figure 9.d), SMOOTHIE can only flatten the landscape to .5*5.6=2.8 (see Figure 9.c, which is still quite steep). On the other hand, if we start near the mean smoothness of our SE data (see Figure 9.b) then a 50% flattening takes us into a quite flat region (see Figure 9.a) where models generalize better.
@@
Figure 9: An illustration of sharp minima forming with higher smoothness. Figure 9d is the upper limit of the sharpness that SMOOTHIE can handle (see Figure 7).
@@ 7.2 Threats to Validity
8 Limitations
The major limitation of this work is the choice of measurement of smoothness, and the use of smoothness as a heuristic. For example, having derived the Hessian, the condition number is an alternate measure of the “nice nature” of the landscape [11].
Another clear limitation is that this approach is limited to learners for which a smoothness can be derived. Especially in the case where the Hessian may not exist (e.g., if the loss function is not twice-differentiable such as the mean absolute error loss), this approach fails. Nevertheless, we have shown successful application in a classical learner that is not gradient-based, Naive Bayes. In that case, we used the negative log-likelihood as a surrogate for the loss function.
A further limitation of this work is that because the smoothness computation is different for each type of learner, this can potentially make the implementation less trivial, especially if the practitioner is interested in trying multiple learners. Our reference implementation uses function overloading to choose the correct equations given the type of the classifier object. The white-box nature of our method means that practitioners need to select a list of classifiers a priori and implement the corresponding smoothness equations.
Finally, we reiterate that SMOOTHIE performs poorly on datasets with high smoothness, as evidenced by our results on the 2-class issue close time prediction results.
9 Future Work
This paper showed that for non-neural learners, we can use the negative log-likelihood as a surrogate for the loss (using Gaussian Naive Bayes). Our future work will expand this to other learners, such as tree-based learners, which have strong performance on tabular data [28]. These algorithms are based on maximizing the information gain (synonymous with the KL divergence) from splitting nodes (which hold subsets of the dataset) on different features. Suppose node $N$ is split into nodes $\{N_{i}\}_{i=1\ldots m}$. Then, the information gain is:
$IG(N)=H(N)-\sum\limits_{i=1}^{m}\frac{\lvert N_{i}\rvert}{\lvert N\rvert}H(N_{i})$
where $H(\cdot)$ is an entropy measure. Most tree-based algorithms use either the Gini index or the Shannon entropy for $H$, but it is worth noting that both are special cases of Tsallis’ entropy [62]. We can then use the negative log-likelihood, which is given by
$\ell:=\text{NLL}=-2\lvert E\rvert\cdot IG$
which follows since the log-likelihood (also called the G-statistic) is a scaled version of the KL divergence. However, since decision trees are non-parametric, we cannot use the Hessian-based smoothness measure; instead, we need to use a center-difference approximation for the smoothness:
$\beta(x)=\frac{\ell(x-h)+\ell(x+h)-2\ell(x)}{h^{2}}$
where $x$ is some point, and $h$ is a small constant. Different methods of estimating the smoothness of the overall learner based on the above can be used, such as a Monte Carlo estimation.